OpenStudy (anonymous):

I am to a certian point and I need to know if I simplify further 3(x+2)(x-9)=(x+7)/3

OpenStudy (anonymous):

or better yet how do I solve the equation when I get to (x+7)/3

OpenStudy (anonymous):

The original equation was this; log3(x+7)-1=log3(x+2)-log3(x-9)

OpenStudy (anonymous):

log3 [ (x + 7)/ 3 ] = log3 [(x+2)/ (x +9)] -> (x + 7)/ 3 = (x+2)/ (x +9) -> (x + 7) (x +9) = 3(x+2) -> x^2 -16x + 63 = 3x + 6 => x^2 - 19x + 57 = 0